You bought a bond five years ago for $856 per bond. The bond is now selling for $850. It also paid $60 in interest per year, which you reinvested in the bond. Calculate the realized rate of return earned on this bond. (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16)) Realized rate of return 631 3196 rch
Answer #1

5 Years ago price =856
Today price =850
Total Interest =60*5 =300
Realized return of this bond =(850-856+300)/856=34.35%